# Learning Monotonic Attention in Transducer for Streaming Generation

### v1

@@ Abstract
Streaming generation models are increasingly utilized across various fields, with the Transducer architecture being particularly popular in industrial applications. However, its input-synchronous decoding mechanism presents challenges in tasks requiring non-monotonic alignments, such as simultaneous translation, leading to suboptimal performance in these contexts. In this research, we address this issue by tightly integrating Transducer’s decoding with the history of input stream via a learnable monotonic attention mechanism. Our approach leverages the forward-backward algorithm to infer the posterior probability of alignments between the predictor states and input timestamps, which is then used to estimate the context representations of monotonic attention in training. This allows Transducer models to adaptively adjust the scope of attention based on their predictions, avoiding the need to enumerate the exponentially large alignment space. Extensive experiments demonstrate that our MonoAttn-Transducer significantly enhances the handling of non-monotonic alignments in streaming generation, offering a robust solution for Transducer-based frameworks to tackle more complex streaming generation tasks. Codes: https://github.com/ictnlp/MonoAttn-Transducer
@@ 1 Introduction
Popular streaming generation methods can be broadly divided into two categories: Attention-based Encoder-Decoder (AED; Bahdanau et al., 2015) and Transducer (Graves, 2012). Streaming AED models adapt the conventional sequence-to-sequence framework (Bahdanau, 2014) to support streaming generation. They often rely on an external policy module to determine the READ/WRITE actions in inference and to direct the scope of cross-attention in training. Examples include Wait-$k$ policy (Ma et al., 2019) and monotonic attention-based methods (Raffel et al., 2017; Arivazhagan et al., 2019; Ma et al., 2020d; 2023a). On the other hand, Transducer models connect the encoder and predictor through a joiner rather than using cross-attention. As shown in Figure 1(a), the joiner is designed to synchronize the encoder and predictor by expanding its output vocabulary to include a blank symbol $\epsilon$, which indicates a READ action. Due to the decoupling of the predictor state from the encoder state, READ/WRITE states in Transducer can be represented by a two-dimensional lattice. This allows for the computation of total probabilities using the forward-backward algorithm (Graves, 2012), facilitating end-to-end optimization. Benefited from joint optimization of all potential policies during training, Transducer often demonstrates better performance compared to AED models (Xue et al., 2022; Wang et al., 2023).
@@
During the decoding process of Transducer, each target token is explicitly aligned with a corresponding source token. This input-synchronous decoding property makes the architecture well-suited for tasks like speech recognition, where the input and output align monotonically. However, it poses challenges for non-monotonic alignment tasks such as simultaneous translation (Chuang et al., 2021; Shao & Feng, 2022; Ma et al., 2023c). Due to the decoupled design, Transducer models have limited ability to attend to the input stream history during decoding, making it hard to manage reorderings. To address this issue, recent research (Liu et al., 2021; Tang et al., 2023) has started to explore the incorporation of cross-attention mechanism to enhance the capacity for handling complex non-monotonicity. Despite these efforts, the integration of cross-attention presents significant challenges. By integrating the predictor states with source history through attention, the representation of predictor states becomes relevant not only to the encoder states but also to the specific READ/WRITE path history (Tang et al., 2023). This results in an exponentially large state space for Transducer, hindering the application of the forward-backward algorithm for end-to-end training.
(a) Transducer
(b) MonoAttn-Transducer (train)
(c) MonoAttn-Transducer (infer)
Figure 1: Illustration of MonoAttn-Transducer. During inference, the predictor state can attend to all generated encoder states through cross-attention. During training, the scope of cross-attention is adjusted based on the posterior alignment $\pi_{u,t}$ derived from the model’s prediction (Equation 5).
@@
In this research, we present an efficient training algorithm for Transducer models to learn the monotonic cross-attention mechanism. This allows Transducer’s predictor to access source history in real-time inference (Figure 1(c)), improving its ability to handle tasks with non-monotonic alignments. As illustrated in Figure 1(b), we leverage the forward-backward algorithm to infer the posterior probability of alignments between predictor and encoder states in training. This derived posterior alignment enables the estimation of context representation for each predictor state using expected soft attention. In this way, Transducer models adaptively adjust the scope of attention based on their predictions, avoiding the need to enumerate the exponentially large alignment space during training.
@@
We conduct experiments on both speech-to-text and speech-to-speech simultaneous translation to demonstrate the generality of our approach across various modalities. MonoAttn-Transducer shows significant improvements in generation quality without a noticeable increase in latency in both ideal and computation-aware settings (§5). Further analysis reveals that MonoAttn-Transducer is particularly effective in handling samples with higher levels of non-monotonicity (§6).
@@ 2.2 Transducer
As shown in Figure 1(a), Transducer model (Graves, 2012) comprises three components: an encoder, a predictor, and a joiner. The encoder unidirectionally encodes the received input prefix ${x}_{1:t}$ into a context representation $h_{t}$. The predictor functions similarly to an autoregressive language model, encoding the dependencies between tokens in the generated prefix ${y}_{1:u}$ into $s_{u}$. The joiner makes predictions based on the current source representation $h_{t}$ and target representation $s_{u}$. If the model needs to READ more information to update the source representation for continued generation, a blank token $\epsilon$ is generated. Otherwise, a WRITE operation is performed, and the generated token is fed back into the predictor to obtain a new target representation. Each time $h_{t}$ or $s_{u}$ is updated, the joiner performs a prediction step until the entire source has been processed. The encoder and predictor are usually modeled using either a recurrent neural network (Graves, 2012) or Transformer layers (Zhang et al., 2020). The joiner is typically composed of a feed-forward network.
@@
$\alpha(t,u)\coloneq\mathrm{Pr}({y}_{1:u}|{x}_{1:t})$ (1)
$\beta(t,u)\coloneq\mathrm{Pr}({y}_{u+1:U}|{x}_{t:T})$
@@
$\alpha(t,u)$ $=\alpha(t-1,u)\mathrm{Pr}(\epsilon|t-1,u)+\alpha(t,u-1)\mathrm{Pr}(y_{u}|t,u-1)$ (2)
$\beta(t,u)$ $=\beta(t+1,u)\mathrm{Pr}(\epsilon|t,u)+\beta(t,u+1)\mathrm{Pr}(y_{u+1}|t,u)$
with initial condition $\alpha(1,0)=1$ and $\beta(T,U)=\mathrm{Pr}(\epsilon|T,U)$. $\mathrm{Pr}(v|t,u)$ denotes the probability of generating token $v$ from $h_{t}$ and $s_{u}$, $v\in\mathcal{V}\cup\{\epsilon\}$. The total output probability is:
$\mathrm{Pr}(\bm{y}|\bm{x})=\alpha(T,U)\mathrm{Pr}(\epsilon|T,U).$ (3)
@@ 3.2.1 Posterior Alignment
Suppose we have a probability lattice $\mathrm{Pr}(v|t,u)$, representing the probability of generating token $v$ from $h_{t}$ and $s_{u}$, for $1\leq t\leq T$, $0\leq u\leq U$, and $v\in\mathcal{V}\cup\{\epsilon\}$. The posterior probability of generating $y_{u}$ at the moment $x_{t}$ is read can be represented by:
$\pi_{u,t}=\dfrac{\mathrm{Pr}({y}_{1:u-1}|{x}_{1:t})\mathrm{Pr}(y_{u}|t,u-1)\mathrm{Pr}({y}_{u+1:U}|{x}_{t:T})}{\mathrm{Pr}({y}_{1:U}|{x}_{1:T})}$ (5)
@@
which implies that the predictor state $s_{0}$ is generated immediately after the first source token arrives. Using the forward and backward variables introduced in Section 2.2, Equation 5 can be concisely expressed as follows:
$\pi_{u,t}=\dfrac{\alpha(t,u-1)\mathrm{Pr}(y_{u}|t,u-1)\beta(t,u)}{\alpha(T,U)\mathrm{Pr}(\epsilon|T,U)}.$ (7)
@@ 3.2.2 Monotonic Attention
This indicates that the expected context representation $c_{u}$ is a weighted sum of context representations under various amount of source information, with the weights given by the posterior alignment probability $\pi_{u,t}$. The nested summation operations in Equation 8 may lead to an increase in computational complexity. Fortunately, Arivazhagan et al. (2019) suggests that it can be rewritten as:
@@
Equation 9 can then be computed efficiently using cumulative sum operations (Arivazhagan et al., 2019).
@@ 3.2.3 Training with Prior Alignment
In the following, we will use $p_{0,t}^{\mathrm{dia}}$ as the default choice for prior alignment and compare the differences between using $p_{0,t}^{\mathrm{uni}}$ and $p_{0,t}^{\mathrm{dia}}$ in the ablation study (Section 6.1).
@@ 3.2.4 Chunk Synchronization
The prior alignment probability is adjusted in the same manner. We detail the entire training process in Algorithm 1.
@@ 4 Related Work
Our work is also related to researches in designing attention modules for streaming AED models. These works often introduce Bernoulli variables to indicate READ/WRITE actions. The distribution of these variables is used to estimate monotonic alignment and to compute the expected context representation in training (Raffel et al., 2017). Depending on the setting of attention window, these works can be classified into monotonic hard attention (Raffel et al., 2017), monotonic chunkwise attention (MoChA; Chiu & Raffel, 2018), and monotonic infinite lookback attention (MILk; Arivazhagan et al., 2019). Ma et al. (2020d) subsequently introduced the MILk mechanism to Transformer models, and Ma et al. (2023b) further proposed a numerically-stable algorithm for estimating monotonic alignment. Unlike the aforementioned works, our approach learns monotonic attention based on the posterior alignment of Transducer, avoiding the use of unstable Bernoulli variables.
@@ 5 Experiments
We validate the performance of our MonoAttn-Transducer on two typical streaming generation tasks: speech-to-text and speech-to-speech simultaneous translation. The differences in grammatical structures between the source and target languages often necessitate word reordering during generating translation. This property makes the simultaneous translation task well-suited for evaluating the ability of MonoAttn-Transducer in handling non-monotonic alignments.
@@ 5.1 Experimental Setup
Evaluation We use SimulEval toolkit (Ma et al., 2020b) for evaluation. Translation quality is assessed using case-sensitive detokenized BLEU (Papineni et al., 2002; Post, 2018) and neural-based COMET-22 score. Latency is measured by word-level Average Lagging (AL; Ma et al., 2019; 2020c). [^1] Numerical results with more metrics are provided in Appendix C. Notably, Table 7 presents a comparison of the computation-aware latency metrics for AL and LAAL (Papi et al., 2022) between the Transducer and MonoAttn-Transducer models. For speech-to-speech experiments, translation quality is assessed using ASR-BLEU and latency is measured by delay of generated waveform chunks (Ma et al., 2022).
Figure 2: An example of diagonal prior and posterior alignment from MuST-C English-to-Spanish training corpus. The vertical axis represents the target subword sequence and the horizontal axis represents the speech waveform. Darker areas indicate higher alignment probabilities. Chunk size in this example is set to 640ms. More examples are provided in Appendix D.
@@ 5.2 Main Results
Moreover, we have observed some notable results of MonoAttn-Transducer-Prior. With a larger chunk size, the performance of MonoAttn-Transducer-Prior is comparable to that of MonoAttn-Transducer-Posterior, and even slightly outperforming the latter in En-De. However, there exists a significant performance drop with a smaller chunk size. Specifically, with a chunk size of 320ms, MonoAttn-Transducer-Prior’s generation quality is on average 1.03 BLEU lower than Transducer baseline under similar latency conditions. This phenomenon highlights the importance of learning monotonic attention through inferring posterior alignment. From the chunk synchronization mechanism described in Equation 12, smaller chunk sizes require finer alignment granularity between the predictor and encoder states. This increased granularity necessitates more precise alignment to estimate the expected context representation during training. Figure 2 provides an example of diagonal prior and posterior alignment. While the diagonal prior generally captures the trend of the alignment information, it can be skewed by the uneven distribution of speech information and possible local reorderings. In contrast, the inferred posterior offers a more confident and accurate alignment probability. For instance, the diagonal prior assigns a high probability to aligning the word “si (if)” with the timestep preceding the waveform of “if”, while the inferred posterior corrects this misalignment. Therefore, learning monotonic attention with posterior alignment leads to a more accurate estimation of context representation and improved performance. [^2] We present a comparison between the prior and posterior under various chunk sizes in Appendix D. A key observation is that as the alignment granularity becomes finer, the differences gradually increases. In subsequent experiments, we represent MonoAttn-Transducer using the results of MonoAttn-Transducer-Posterior.
@@ 5.3 Comparison with State-of-the-Art
We compare MonoAttn-Transducer with state-of-the-art open-source approaches in simultaneous translation, including Wait-$k$ (Ma et al., 2020c), RealTrans (Zeng et al., 2021), CAAT (Liu et al., 2021), MU-ST (Zhang et al., 2022), EDAtt (Papi et al., 2023), Seg2Seg (Zhang & Feng, 2023) and NAST (Ma et al., 2024). Further details about baselines are available in Appendix B. Results are plotted in Figure 3(a) and 3(b). We observe that learning monotonic attention significantly enhances the performance of Transducer, making it comparable to state-of-the-art models. Compared to CAAT, another Transducer-based model, MonoAttn-Transducer demonstrates superiority in scenarios with less stringent latency requirements. Under a latency of approximately 2s, it outperforms CAAT by 1.1 BLEU in En-De. This clearly demonstrates the advantage of MonoAttn-Transducer’s tightly coupled self-attention and cross-attention modules in the predictor, which facilitates the learning of richer representations.
As discussed in Section 4, TAED is another Transducer-based model highly relevant to our work. However, the code and distilled data used to train TAED in Tang et al. (2023) have not been made publicly available. This lack of open access hinders a fair comparison of TAED with our MonoAttn-Transducer. Despite this, we attempt to analyze the performance by comparing each with Transducer baseline in their respective experimental settings. The comparison is shown in Table 8. We have observed that the improvement from TAED is more pronounced with smaller chunk sizes, which contrasts with the results of MonoAttn-Transducer. We speculate that this is because, in TAED, the representations of all generated predictor states are updated every time the encoder receives a new speech chunk. This helps TAED generate more accurate representations when the chunk size is small. However, this mechanism in TAED incurs an $O(T+U)$ forward propagation cost during simultaneous inference, which can significantly increase latency in practice due to heavy computational overhead when the chunk size is small. In contrast, MonoAttn-Transducer maintains an $O(U)$ complexity as Transducer baseline. As shown in Table 7, this property minimizes the gap between ideal and computation-aware latency, offering advantages in real-time applications.

### v2

# Overcoming Non-monotonicity in Transducer-based Streaming Generation

@@ Abstract
Streaming generation models are utilized across fields, with the Transducer architecture being popular in industrial applications. However, its input-synchronous decoding mechanism presents challenges in tasks requiring non-monotonic alignments, such as simultaneous translation. In this research, we address this issue by integrating Transducer’s decoding with the history of input stream via a learnable monotonic attention. Our approach leverages the forward-backward algorithm to infer the posterior probability of alignments between the predictor states and input timestamps, which is then used to estimate the monotonic context representations, thereby avoiding the need to enumerate the exponentially large alignment space during training. Extensive experiments show that our MonoAttn-Transducer effectively handles non-monotonic alignments in streaming scenarios, offering a robust solution for complex generation tasks. Code is available at https://github.com/ictnlp/MonoAttn-Transducer.
Machine Learning, ICML
@@ 1 Introduction
Popular streaming generation methods can be broadly divided into two categories: Attention-based Encoder-Decoder (AED; Bahdanau et al., 2015) and Transducer (Graves, 2012). Streaming AED models adapt the conventional sequence-to-sequence framework (Bahdanau, 2014) to support streaming generation. They often rely on an external policy module to determine the READ/WRITE actions in inference and to direct the scope of cross-attention in training. Examples include Wait-$k$ policy (Ma et al., 2019) and monotonic attention-based methods (Raffel et al., 2017; Arivazhagan et al., 2019; Ma et al., 2020d, 2023a). On the other hand, Transducer models connect the encoder and predictor through a joiner rather than using cross-attention. The joiner is designed to synchronize the encoder and predictor by expanding its output vocabulary to include a blank symbol $\epsilon$, which indicates a READ action. Due to the decoupling of the predictor state from the encoder state, READ/WRITE states in Transducer can be represented by a two-dimensional lattice. This allows for the computation of total probabilities using the forward-backward algorithm (Graves, 2012), facilitating end-to-end optimization. Benefited from joint optimization of all potential policies during training, Transducer often demonstrates better performance compared to AED models (Xue et al., 2022; Wang et al., 2023).
@@
During the decoding process of Transducer, each target token is explicitly aligned with a corresponding source token. This input-synchronous decoding property makes the architecture well-suited for tasks like speech recognition, where the input and output align monotonically. However, it poses challenges for non-monotonic alignment tasks such as simultaneous translation (Chuang et al., 2021; Shao & Feng, 2022; Ma et al., 2023b, c). Due to the decoupled design, Transducer models have limited ability to attend to the input stream history during decoding, making it hard to manage reorderings. To address this issue, recent research (Liu et al., 2021; Tang et al., 2023) has started to explore the incorporation of cross-attention mechanism to enhance the capacity for handling complex non-monotonicity. Despite these efforts, the integration of cross-attention presents significant challenges. By integrating the predictor states with source history through attention, the representation of predictor states becomes relevant not only to the encoder states but also to the specific READ/WRITE path history (Tang et al., 2023). This results in an exponentially large state space for Transducer, hindering the application of the forward-backward algorithm for end-to-end training.
@@
In this research, we present an efficient training algorithm for Transducer models to learn the monotonic cross-attention mechanism. This allows Transducer’s predictor to access source history in real-time inference, improving its ability to handle tasks with non-monotonic alignments. We leverage the forward-backward algorithm to infer the posterior probability of alignments between predictor and encoder states in training. This derived posterior alignment enables the estimation of context representation for each predictor state using expected soft attention. In this way, Transducer models adaptively adjust the scope of attention based on their predictions, avoiding the need to enumerate the exponentially large alignment space during training.
@@
We conduct experiments on both speech-to-text/speech simultaneous translation to demonstrate the generality of our approach across various modalities. MonoAttn-Transducer shows significant improvements in generation quality without a noticeable increase in latency in both ideal and computation-aware settings (§5). Further analysis reveals that MonoAttn-Transducer is particularly effective in handling samples with higher levels of non-monotonicity (§6).
@@ 2.2 Transducer
Transducer model (Graves, 2012) comprises three components: an encoder, a predictor, and a joiner. The encoder unidirectionally encodes the received input prefix ${x}_{1:t}$ into a context representation $h_{t}$. The predictor functions similarly to an autoregressive language model, encoding the dependencies between tokens in the generated prefix ${y}_{1:u}$ into $s_{u}$. The joiner makes predictions based on the current source representation $h_{t}$ and target representation $s_{u}$. If the model needs to READ more information to update the source representation for continued generation, a blank token $\epsilon$ is generated. Otherwise, a WRITE operation is performed, and the generated token is fed back into the predictor to obtain a new target representation. Each time $h_{t}$ or $s_{u}$ is updated, the joiner performs a prediction step until the entire source has been processed. The encoder and predictor are usually modeled using either a recurrent neural network (Graves, 2012) or Transformer layers (Zhang et al., 2020). The joiner is typically composed of a feed-forward network.
@@
$\alpha(t,u)\coloneq p({y}_{1:u}|{x}_{1:t})$ (1)
$\beta(t,u)\coloneq p({y}_{u+1:U}|{x}_{t:T})$
@@
$\alpha$ $(t,u)$ (2)
$=\alpha(t-1,u)p(\epsilon|t-1,u)+\alpha(t,u-1)p(y_{u}|t,u-1)$
$\beta$ $(t,u)$
$=\beta(t+1,u)p(\epsilon|t,u)+\beta(t,u+1)p(y_{u+1}|t,u)$
with initial condition $\alpha(1,0)=1$ and $\beta(T,U)=p(\epsilon|T,U)$. $p(v|t,u)$ denotes the probability of generating token $v$ from $h_{t}$ and $s_{u}$, $v\in\mathcal{V}\cup\{\epsilon\}$. The total output probability is:
$p(\bm{y}|\bm{x})=\alpha(T,U)p(\epsilon|T,U).$ (3)
@@ 3 Method
In this section, we provide a detailed introduction to our proposed MonoAttn-Transducer.
@@ 3.2.1 Posterior Alignment
Suppose we have a probability lattice $p(v|t,u)$, representing the probability of generating token $v$ from $h_{t}$ and $s_{u}$, for $1\leq t\leq T$, $0\leq u\leq U$, and $v\in\mathcal{V}\cup\{\epsilon\}$. The posterior probability of generating $y_{u}$ at the moment $x_{t}$ is read can be represented by:
$\pi_{u,t}=\dfrac{p({y}_{1:u-1}|{x}_{1:t})p(y_{u}|t,u-1)p({y}_{u+1:U}|{x}_{t:T})}{p({y}_{1:U}|{x}_{1:T})}$ (5)
@@
which implies that the predictor state $s_{0}$ is generated immediately after the first source token arrives. Using the forward and backward variables introduced in Section 2.2, Eq. 5 can be concisely expressed as follows:
$\pi_{u,t}=\dfrac{\alpha(t,u-1)p(y_{u}|t,u-1)\beta(t,u)}{\alpha(T,U)p(\epsilon|T,U)}.$ (7)
@@ 3.2.2 Monotonic Attention
This indicates that the expected context representation $c_{u}$ is a weighted sum of context representations under various amount of source information, with the weights given by the posterior alignment probability $\pi_{u,t}$. The nested summation operations in Eq. 8 may lead to an increase in computational complexity. Fortunately, Arivazhagan et al. (2019) suggests that it can be rewritten as:
@@
Eq. 9 can then be computed efficiently using cumulative sum operations (Arivazhagan et al., 2019). Then in training, Eq. 4 can be estimated as
$s_{u}=f_{\theta}(s_{0:u-1},c_{u}).$ (10)
@@ 3.2.4 Chunk Synchronization
The prior alignment probability is adjusted in the same manner. We detail the entire training process in Alg. 1.
@@ 4 Related Work
Our work is also related to researches in designing attention modules for streaming AED models. These works often introduce Bernoulli variables to indicate READ/WRITE actions. The distribution of these variables is used to estimate monotonic alignment and to compute the expected context representation in training (Raffel et al., 2017). Depending on the setting of attention window, these works can be classified into monotonic hard attention (Raffel et al., 2017), monotonic chunkwise attention (MoChA; Chiu & Raffel, 2018), and monotonic infinite lookback attention (MILk; Arivazhagan et al., 2019). Ma et al. (2020d) subsequently introduced the MILk mechanism to Transformer models, and Ma et al. (2023a) further proposed a numerically-stable algorithm for estimating monotonic alignment. Unlike the aforementioned works, our approach learns monotonic attention based on the posterior alignment of Transducer, avoiding the use of unstable Bernoulli variables.
@@ 5 Experiments
We validate the performance of our MonoAttn-Transducer on two typical streaming generation tasks: speech-to-text and speech-to-speech simultaneous translation. The differences in grammatical structures between the source and target languages often necessitate word reordering during generating translation. This property makes the simultaneous translation task well-suited for evaluating the ability in handling non-monotonic alignments.
@@ 5.1 Experimental Setup
Evaluation We use SimulEval toolkit (Ma et al., 2020b) for evaluation. Translation quality is assessed using case-sensitive detokenized BLEU (Papineni et al., 2002; Post, 2018) and neural-based COMET-22 score. Latency is measured by word-level Average Lagging (AL; Ma et al., 2019, 2020c) and Length-Adaptive Average Lagging (LAAL; Papi et al., 2022). [^1] Numerical results with more metrics are provided in App. C. Notably, Table 6 presents a comparison of the computation-aware latency metrics for AL and LAAL between the Transducer and MonoAttn-Transducer models. For speech-to-speech tasks, translation quality is assessed using ASR-BLEU and latency is measured by delay of generated waveform chunks (Ma et al., 2022).
Figure 1: An example of diagonal prior and posterior alignment from MuST-C English-to-Spanish training corpus. The vertical axis represents the target subword sequence and the horizontal axis represents the speech waveform. Darker areas indicate higher alignment probabilities. Chunk size in this example is set to 640ms. More examples are provided in App. D.
@@ 5.2 Main Results
Moreover, we have observed some notable results of MonoAttn-Transducer-Prior. With a larger chunk size, the performance of MonoAttn-Transducer-Prior is comparable to that of MonoAttn-Transducer-Posterior, and even slightly outperforming the latter in En-De. However, there exists a significant performance drop with a smaller chunk size. Specifically, with a chunk size of 320ms, MonoAttn-Transducer-Prior’s generation quality is on average 1.03 BLEU lower than Transducer baseline under similar latency conditions. This phenomenon highlights the importance of learning monotonic attention through inferring posterior alignment. From the chunk synchronization mechanism described in Eq. 13, smaller chunk sizes require finer alignment granularity between the predictor and encoder states. This increased granularity necessitates more precise alignment to estimate the expected context representation during training. Figure 1 provides an example of diagonal prior and posterior alignment. While the diagonal prior generally captures the trend of the alignment information, it can be skewed by the uneven distribution of speech information and possible local reorderings. In contrast, the inferred posterior offers a more confident and accurate alignment probability. For instance, the diagonal prior assigns a high probability to aligning the word “si (if)” with the timestep preceding the waveform of “if”, while the inferred posterior corrects this misalignment. Therefore, learning monotonic attention with posterior alignment leads to a more accurate estimation of context representation and improved performance. [^2] We present a comparison between the prior and posterior under various chunk sizes in App. D. A key observation is that as the alignment granularity becomes finer, the differences gradually increases. In subsequent experiments, we represent MonoAttn-Transducer using the results of MonoAttn-Transducer-Posterior.
@@ 5.3 Comparison with State-of-the-Art
We compare MonoAttn-Transducer with state-of-the-art open-source approaches in simultaneous translation, including Wait-$k$ (Ma et al., 2020c), RealTrans (Zeng et al., 2021), CAAT (Liu et al., 2021), MU-ST (Zhang et al., 2022), EDAtt (Papi et al., 2023a), AlignAtt (Papi et al., 2023b), Seg2Seg (Zhang & Feng, 2023), NAST (Ma et al., 2024) and LLM-based LSG (Guo et al., 2025). Further details about baselines are available in App. B. Results are plotted in Figure 2(a) and 2(b). We observe that learning monotonic attention significantly enhances the performance of Transducer, achieving leading streaming generation quality. Compared to CAAT, another Transducer-based model, MonoAttn-Transducer demonstrates superiority in scenarios with less stringent latency requirements. This clearly demonstrates the advantage of MonoAttn-Transducer’s tightly coupled self-attention and cross-attention modules in the predictor, which facilitates the learning of richer representations.
TAED is another Transducer-based model highly relevant to our work. However, the code and distilled data used to train TAED in Tang et al. (2023) have not been made publicly available. This lack of open access hinders a fair comparison of TAED with our MonoAttn-Transducer. Despite this, we attempt to analyze the performance by comparing each with Transducer baseline in their respective experimental settings. The comparison is shown in Table 7 in App. C. We have observed that the improvement from TAED is more pronounced with smaller chunk sizes, which contrasts with the results of MonoAttn-Transducer. We speculate that this is because, in TAED, the representations of all generated predictor states are updated every time the encoder receives a new speech chunk. This helps TAED generate more accurate representations when the chunk size is small. However, this mechanism in TAED incurs an $O(T+U)$ forward propagation cost during simultaneous inference, which can significantly increase latency in practice due to heavy computational overhead when the chunk size is small. In contrast, MonoAttn-Transducer maintains an $O(U)$ complexity as Transducer baseline. As shown in Table 6, this property minimizes the gap between ideal and computation-aware latency, offering advantages in real-time applications.